Translate the following situation into an inequality. Do not solve.
“Marla plans to go on a short vacation and has budgeted $45 dollars per day for meals and the hotel is $165 per day. How many days she can stay, if she has $750 to spend on hotel and meals?”

Answer by stanbon(75887)

Marla plans to go on a short vacation and has budgeted $45 dollars per day for meals and the hotel is $165 per day. How many days she can stay, if she has $750 to spend on hotel and meals?”
----------------------
Daily cost: 45 + 165 = 210
----
# of days = 750/210 = 3.57 days
--------------------------------------
Another way to look at it.
750/165 = 4.55
---
She could stay 4 days at the hotel and have $90 left for meals.
---
90/12 = $7.50 for each of 12 meals over 4 days.
==========================
It's your choice.
==========================
Cheers,
Stan H.